# When Collaborative Maintenance Falls Short: The Persistence of Retracted Papers on Wikipedia

### v2

@@ Abstract
Wikipedia serves as a key infrastructure for public access to scientific knowledge, but it faces challenges in maintaining the credibility of cited sources—especially when scientific papers are retracted. This paper investigates how citations to retracted research are handled on English Wikipedia. We construct a novel dataset that integrates Wikipedia revision histories with metadata from Retraction Watch, Crossref, Altmetric, and OpenAlex, identifying 1,181 citations of retracted papers. We find that 71.6% of all citations analyzed are problematic. These are citations added before a paper’s retraction, as well as the citations introduced after retraction without any in-text mention of the paper’s retracted status. Our analysis reveals that these citations persist for a median of over 3.68 years (1,344 days). Through survival analysis, we find that signals of human attention are associated with a faster correction process. Unfortunately, a paper’s established scholarly authority—a higher academic citation count—is associated with a slower time to correction. Our findings highlight how the Wikipedia community supports collaborative maintenance but leaves gaps in citation-level repair. We contribute to CSCW research by advancing our understanding of this sociotechnical vulnerability, which takes the form of a community coordination challenge, and by offering design directions to support citation credibility at scale.
@@ 1 Introduction
Wikipedia is one of the most influential collaborative knowledge infrastructures in the world. Maintained by a decentralized network of human contributors, automated bots, and community governance processes, it produces and maintains a vast, continuously evolving corpus of public knowledge forte2013defining ; halfaker2012bots ; geiger2010work . This content shapes public understanding across scientific, cultural, and political domains. It is increasingly used not only by students and journalists but also by search engines and large language models as a foundational information source touvron2023llama ; brown2020language . Given its centrality in the modern knowledge ecosystem, the ability of the Wikipedia community to maintain the integrity of its information over time has profound epistemic and societal consequences fallis2008toward ; aragon2021preliminary .
@@
A growing body of CSCW and HCI research has examined how the Wikipedia community responds to real-world change, particularly in the context of high-attention events such as natural disasters, political controversies, social movements, and celebrities’ deaths keegan2013hot ; keegan2012staying ; twyman2017blm ; keegan2015is . These studies have shown that Wikipedia can mobilize contributors to rapidly produce and update information in response to new developments, demonstrating the community’s capacity for collaborative sensemaking. However, far less is known about how well Wikipedia’s editors and community processes handle low-salience epistemic disruptions like paper retractions—cases where previously accepted knowledge becomes invalidated after its initial citation on the platform. These disruptions are slower, less visible, and often go unnoticed by editors and readers alike. The persistence of such trackable errors demonstrates a fundamental coordination challenge within large-scale, decentralized communities: how to ensure that critical updates are routed to the specific locations where repair is needed kittur2008harnessing ; forte2012coordination .
@@
Figure 1: Estimated cumulative and monthly page views of English Wikipedia articles that cite retracted papers without in-text mention of retraction status. The figure only shows citations with no accompanying retraction-related keywords (“retract”, “withdraw”, or “invalid”) in the main text. This keyword-based detection of uncorrected retraction provides a coarser approximation than the correction categorization used in the main analysis.
@@
Our study focuses on scientific retractions, one of the clearest signals of post-publication epistemic breakdown jackson201411 ; fallis2008toward . Retractions occur when scientific findings are withdrawn due to error, misconduct, or other credibility issues, and they function as formal invalidations of prior knowledge fang2012misconduct . When such papers remain cited on Wikipedia without mention of retraction status or removal, they present a critical test of the community’s ability to maintain trustworthy knowledge. From the perspective of readers and other downstream users, this poses serious risks. A person seeking information about a medical treatment, scientific finding, or public health claim may unknowingly encounter content on Wikipedia grounded in invalidated evidence, without any clear warning or correction. Such lapses not only risk misinforming individuals in critical decision-making contexts but also erode trust in Wikipedia as a reliable source of up-to-date knowledge.
@@
As is illustrated in Figure 1, English Wikipedia articles containing uncorrected citations to retracted papers have received over 250 million cumulative page views as of 10/24/2024. While not every view results in exposure to the specific citation or misinformation, this figure reflects the substantial reach of flawed content. Moreover, given Wikipedia’s role as a foundational information layer across the web, including in search engine results, educational settings, and machine learning training data, the consequences of uncorrected retractions extend far beyond individual pages and would be difficult to quantify thelwall2025does .
@@
Rather than analyzing editor behavior directly, this paper takes a system-level perspective on how the Wikipedia community handles retracted scientific knowledge. Building on CSCW scholarship on collaborative infrastructures, knowledge governance, and epistemic maintenance ackerman2013sharing ; aragon2021preliminary ; forte2012coordination , we examine how citations of retracted papers are distributed across the platform and what observable features are associated with how quickly these citations are corrected. To do so, our study focuses on corrective actions that are directly visible to a typical reader within the main body of an article. We therefore define a “correction” narrowly as either the complete removal of the citation or the addition of a visible in-text warning, which includes explicit mentions of the retraction or editorial notes indicating the need for better sources. This definition intentionally excludes automated templates added only to the reference section, which may not be visible to readers who do not consult the bibliography. This distinction is crucial because it recognizes the division of labor within the Wikipedia community. Automated tools are essential for the scalable work of detection and flagging halfaker2013rise ; halfaker2012bots , providing a vital signal primarily for an editorial audience. However, a complete, reader-facing repair often requires a subsequent act of contextual judgment and editing by humans. Since this human oversight is a key variable and not always guaranteed kharazian2024governance ; baigutanova2023longitudinal , our study’s focus on in-text corrections is designed to measure the success of this complete human-in-the-loop process.
@@
Here, we combine data from Retraction Watch oransky2012retraction , Crossref Events hendricks2020crossref , Altmetric LLC altmetric , OpenAlex priem2022openalex , and Wikipedia revision and page view histories to investigate two core aspects of retracted research on Wikipedia:
@@
  • Overview of retracted paper citations on the English Wikipedia. We explore the distribution of uncorrected citations to retracted papers relative to non-retracted ones and identify the academic domains most susceptible to citing retracted papers on Wikipedia. This descriptive overview quantifies the scope of the issue and its potential impact on English Wikipedia users.
@@
  • What factors influence the speed at which retracted paper citations on English Wikipedia are corrected? Our study investigates the characteristics associated with citations of retracted papers being corrected on Wikipedia. We analyze various features, including a range of features related to the paper itself (e.g., its academic domain, open access status, and academic citation count). Additionally, we examine factors related to the retraction event (e.g. the reasons for retraction, time from publication to retraction, and editorial events before retraction). Finally, we analyzed Wikipedia-specific factors such as the level of editorial activity (e.g., the number of unique editors). These factors provide insight into the challenges of identifying and addressing problematic citations.
Our findings show that Wikipedia’s responsiveness to citations of retracted papers is inconsistent and often slow. While some citations are corrected quickly, many persist uncorrected for years. Among the 1,181 citation pairs analyzed, we found 71.6% to be problematic—either added before the retraction or introduced after the retraction without in-text retraction warnings. Citations are corrected more quickly when the retraction was for serious issues like fraud or error, when Wikipedia editors flagged the paper as controversial at the time of its initial citation, when the paper is open access, and when it received substantial attention on Twitter/X. Corrections are also faster on pages with more unique editors and for papers in the Physical and Social Sciences. Conversely, citations to papers with higher academic citation counts are corrected more slowly, suggesting that a paper’s established impact may create editorial inertia. These patterns demonstrate the complex interplay of visibility, editorial capacity, and contextual salience in Wikipedia’s citation maintenance practices.
@@
Our paper contributes to CSCW research on collaborative knowledge infrastructures in four ways. First, it shifts the focus from Wikipedia’s well-documented responsiveness to high-salience events to its underexamined handling of low-visibility epistemic breakdowns, such as retracted scientific papers. Second, it provides a large-scale empirical analysis of how retracted citations persist—or are corrected—across academic domains, revealing how features such as social media signals, retraction reasons, and editor numbers are associated with repair outcomes. Third, it highlights a disconnect between reference-level flagging and substantive in-text repair. By focusing our analysis on corrections visible within the article body, we highlight how automated signals in the reference section may not translate into the human editorial work required to address problematic citations for a general reader appropriately. This indicates a broader design gap between machine-generated alerts and editorial workflows. Finally, it provides actionable opportunities for design, suggesting how Wikipedia’s tools, workflows, or interfaces might better support editors in identifying and repairing citation-level credibility failures.
@@ 2.1 Wikipedia as an Epistemic Infrastructure
Wikipedia has served as a prime example in CSCW and HCI research for studying large-scale, decentralized collaboration. As a social computing system, it depends on the interplay of human contributors, automated bots, and governance policies to produce and maintain public knowledge forte2013defining ; halfaker2012bots ; geiger2010work . This sociotechnical infrastructure supports a complex ecosystem of roles, routines, and technical tools that enable collaborative knowledge production ackerman2013sharing ; forte2012coordination .
@@
Prior work has demonstrated how Wikipedia adapts to rapidly changing external events. In crisis contexts such as natural disasters or breaking news, editors engage in high-tempo coordination, rapidly creating and revising articles to reflect emerging realities keegan2013hot ; keegan2012staying . Studies of social movement coverage, such as the documentation of Black Lives Matter, extend this perspective to the formation of collective memory in contentious and evolving domains twyman2017blm . Keegan et al. keegan2015is show how editors update Wikipedia biographies after someone dies, working together to correct facts while also honoring the person’s memory. Together, these studies demonstrate Wikipedia’s potential for responsiveness, focusing on high-salience events that trigger rapid editorial attention.
@@
Yet recent work highlights Wikipedia’s vulnerabilities in less visible contexts. Kharazian et al. kharazian2024governance show how, in smaller language editions of Wikipedia, editorial control can become concentrated in the hands of a few users, allowing misinformation to persist. Similarly, Aragón et al. aragon2021preliminary identify risks to knowledge integrity that can arise even without deliberate manipulation. Together, these studies show that Wikipedia’s openness—while foundational to its success—also means that no individual contributor is accountable for the accuracy of any specific claim. When responsibility is diffuse and editorial attention is uneven, epistemic breakdowns can persist without being noticed or repaired. Moreover, Wikipedia’s information credibility can face intentional challenges. Kumar et al. kumar2016disinformation studied the creation and detection of Wikipedia hoaxes, fabricated articles designed to deceive readers. Our study contributes to the literature by examining how Wikipedia handles a particular form of non-adversarial epistemic disruption: the citation of scientific articles that have been formally retracted. Failure to maintain the credibility of cited scientific sources leads to a non-negligible breakdown in information quality that is impactful regardless of intention.
@@
Prior work has provided important insights into Wikipedia’s collaborative maintenance processes at the article level. However, fewer studies have examined how the platform manages the credibility of cited knowledge when it becomes formally invalidated. Rather than tracing editorial workflows directly, we analyze the distribution of these citations across Wikipedia and identify factors associated with whether they are eventually corrected or persist unaddressed. This approach highlights how certain infrastructural and visibility features—such as the number of unique human editors, Twitter/X attention to the papers, and the papers’ academic domains—shape citation repair in decentralized collaborative systems.
@@ 2.2 Scientific Knowledge and Citations on Wikipedia
Scientific knowledge plays a central role in Wikipedia’s coverage of specialized topics, and recent scholarship has explored how scientific sources are cited, structured, and maintained. Prior work has mapped the landscape of scientific citations on Wikipedia, revealing both the breadth of disciplines represented and the dominant role of biomedical research and papers from high-impact journals. Singh, West, and Colavizza singh2021wikipedia quantified this trend, finding that the life sciences and biomedicine dominate the cited literature, with journals such as Nature and Science among the most frequently referenced. Arroyo-Machado et al. arroyo2020science similarly found that multidisciplinary and biomedical journals dominate Wikipedia’s scientific references. Nicholson et al. nicholson2021measuring assessed citation quality, showing that Wikipedia tends to cite reputable journals, but still includes a notable fraction of sources with limited academic validation. Other studies have documented representational biases, such as the under-citation of women and non-Anglophone authors zheng2023gender .
@@
Complementing this citation-level research, scholars have also examined how Wikipedia’s article quality ratings relate to citation behavior. Arroyo-Machado et al. arroyo2022wikinformetrics found that articles rated as higher quality by the Wikipedia community tend to include more academic references, attract more edits, and receive higher page views—but that these associations vary depending on the topical domain. For example, the authors found that a high number of academic references strongly correlates with scholarly-oriented articles in fields like medicine and history. In contrast, high page views are more typical of articles about popular culture, and high talk page activity is associated with controversial subjects. More recently, Baigutanova et al. baigutanova2023longitudinal conducted a large-scale longitudinal study of Wikipedia’s reference quality, finding that sourcing practices have generally improved over time, particularly with increasing reliance on peer-reviewed journals and scientific literature. This trend toward higher-quality sourcing builds on earlier observations about Wikipedia’s increasing preference for high-impact journals such as Nature and Science nielsen2007scientific ; wedemeyer2008quality . Nevertheless, concerns persist about the reliability and maintenance of references, particularly in articles that receive limited editorial attention.
@@
While prior research has primarily focused on the inclusion and quality of citations, relatively little is known about how Wikipedia handles erroneous scientific citations, particularly those referencing retracted research. The study most closely aligned with ours is by Nicholson et al. nicholson2021measuring , who identified 118 retracted articles cited across 297 Wikipedia pages. In a manual sample of 50 citations, they found that about half failed to acknowledge the retraction, and even when retractions were mentioned, they were typically used illustratively rather than corrected or flagged clearly. Additionally, their study was based on a small sample. It did not examine what factors drive citation correction, thus creating an opportunity for a large-scale characterization of retraction persistence patterns across Wikipedia.
@@
We build on this line of work by identifying retracted papers cited on Wikipedia and analyzing how those citations are handled over time through article revisions. Our study offers a large-scale account of how retracted citations are noted, ignored, or removed. We also model the factors that shape these outcomes. In doing so, we extend prior research on citation practices by shifting attention from how scientific citations are sourced and evaluated to how they become invalid and are subsequently addressed. Our work also contributes to CSCW scholarship on collaborative maintenance and infrastructure resilience by examining how Wikipedia editors manage epistemic failures at the citation level.

### v3

# The Persistence of Retracted Papers on Wikipedia

@@ Abstract
Wikipedia serves as a key infrastructure for public access to scientific knowledge, but it faces challenges in maintaining the credibility of cited sources—especially when scientific papers are retracted. This paper investigates how citations to retracted research are handled on English Wikipedia. We construct a novel dataset that integrates Wikipedia revision histories with metadata from Retraction Watch, Crossref, Altmetric, and OpenAlex, identifying 1,181 citations of retracted papers. We find that 71.6% of the citations were initially problematic and in need of reader-facing repair, defined as those added before the paper’s retraction (51.5%) or introduced afterwards without proper warning (20.1%). While many are eventually corrected, our analysis reveals that these citations persist for a median of 3.68 years (1,344 days). Through survival analysis, we find that bot-mediated flagging (RetractionBot), open access availability, pre-existing online visibility (e.g., Twitter/X mention counts), and page-level organization (e.g., number of categories on a Wikipedia page) are associated with a higher hazard of correction. Conversely, a paper’s established scholarly authority—a higher academic citation count—is associated with a slower time to correction. Our findings highlight how the Wikipedia community supports collaborative maintenance but leaves gaps in citation-level repair. We contribute to CSCW research by advancing our understanding of this sociotechnical vulnerability, which takes the form of a community coordination challenge, and by offering design directions to support citation credibility at scale.
@@ 1 Introduction
Wikipedia is one of the most influential collaborative knowledge infrastructures in the world. Maintained by a decentralized network of human contributors, automated bots, and community governance processes, it produces and maintains a vast, continuously evolving corpus of public knowledge Forte and Lampe (2013); Halfaker and Riedl (2012); Geiger and Ribes (2010). This content shapes public understanding across scientific, cultural, and political domains. It is increasingly used not only by students and journalists but also by search engines and large language models as a foundational information source Touvron et al. (2023); Brown et al. (2020). Given its centrality in the modern knowledge ecosystem, the ability of the Wikipedia community to maintain the integrity of its information over time has profound epistemic and societal consequences Fallis (2008); Aragón and Sáez-Trumper (2021).
@@
A growing body of CSCW and HCI research has examined how the Wikipedia community responds to real-world change, particularly in the context of high-attention events such as natural disasters, political controversies, social movements, and celebrities’ deaths Roberts and Xiong-Gum (2022); Kurek et al. (2025); Keegan et al. (2013, 2012); Twyman et al. (2017); Keegan and Brubaker (2015). These studies have shown that Wikipedia can mobilize contributors to rapidly produce and update information in response to new developments, demonstrating the community’s capacity for collaborative sensemaking. However, far less is known about how well Wikipedia’s editors and community processes handle low-salience epistemic disruptions like paper retractions—cases where previously accepted knowledge becomes invalidated after its initial citation on the platform. These disruptions are slower, less visible, and often go unnoticed by editors and readers alike. The persistence of such trackable errors highlights a fundamental coordination challenge within large-scale, decentralized communities: how to ensure that critical updates are routed to the specific locations where repair is needed Kittur and Kraut (2008); Forte et al. (2012).
@@
Figure 1: Estimated cumulative and monthly page views of English Wikipedia articles that cite retracted papers without in-text mention of retraction status. Page view data were retrieved using the Wikimedia REST API based on the citation dataset detailed in Section 3.1. The figure only shows citations with no accompanying retraction-related keywords (“retract”, “withdraw”, or “invalid”) in the main text. This keyword-based detection of uncorrected retraction provides a coarser approximation than the correction categorization used in the main analysis (see Section 3.1 for details on data collection).
@@
Our study focuses on scientific retractions, one of the clearest signals of post-publication epistemic breakdown Jackson (2014); Fallis (2008). Retractions occur when scientific findings are withdrawn due to error, misconduct, or other credibility issues, and they function as formal invalidations of prior knowledge Fang et al. (2012). When such papers remain cited on Wikipedia without mention of retraction status or removal, they present a critical test of the community’s ability to maintain trustworthy knowledge. From the perspective of readers and other downstream users, this poses serious risks. A person seeking information about a medical treatment, scientific finding, or public health claim may unknowingly encounter content on Wikipedia grounded in invalidated evidence, without any clear warning or correction. Such lapses not only risk misinforming individuals in critical decision-making contexts but also erode trust in Wikipedia as a reliable source of up-to-date knowledge.
@@
As is illustrated in Figure 1, English Wikipedia articles containing uncorrected citations to retracted papers have received over 250 million cumulative page views as of 10/24/2024. While not every view results in exposure to the specific citation or misinformation, this figure reflects the substantial reach of flawed content. Moreover, given Wikipedia’s role as a foundational information layer across the web, including in search engine results, educational settings, and machine learning training data, the consequences of uncorrected retractions extend far beyond individual pages and would be difficult to quantify Thelwall et al. (2025).
@@
Rather than analyzing editor behavior directly, this paper takes a system-level perspective on how the Wikipedia community handles retracted scientific knowledge. Building on CSCW scholarship on collaborative infrastructures, knowledge governance, and epistemic maintenance Ackerman et al. (2013); Aragón and Sáez-Trumper (2021); Forte et al. (2012), we examine how citations of retracted papers are distributed across the platform and what observable features are associated with how quickly these citations are corrected. To do so, our study focuses on corrective actions that are directly visible to a typical reader within the main body of an article. We therefore define a “correction” narrowly as either the complete removal of the citation or the addition of a visible in-text warning, which includes explicit mentions of the retraction or editorial notes indicating the need for better sources. This definition intentionally excludes automated templates added only to the reference section, which may not be visible to readers who do not consult the bibliography. At the same time, we treat RetractionBot flagging as an observable feature that may be associated with subsequent reader-facing repair. This distinction is important because it recognizes the division of labor within the Wikipedia community. Automated tools are essential for the scalable work of detection and flagging Halfaker et al. (2013); Halfaker and Riedl (2012), providing a vital signal primarily for an editorial audience. However, a complete, reader-facing repair often requires a subsequent act of contextual judgment and editing by humans. Since this human oversight is a key variable and not always guaranteed Kharazian et al. (2024); Baigutanova et al. (2023), our study’s focus on in-text corrections is designed to measure the success of this complete human-in-the-loop process.
@@
Here, we combine data from Retraction Watch Oransky and Marcus (2012), Crossref Events Hendricks et al. (2020), Altmetric LLC Altmetric Support (2020), OpenAlex Priem et al. (2022), and Wikipedia revision and page view histories to investigate two core aspects of retracted research on Wikipedia:
@@
  • Overview of retracted paper citations on the English Wikipedia. We explore the distribution of uncorrected citations to retracted papers relative to non-retracted ones and identify the academic domains that disproportionately cite retracted papers. This descriptive overview quantifies the scope of the issue and its potential impact on English Wikipedia users.
@@
  • What factors are associated with the timing of corrective actions for retracted paper citations on English Wikipedia? This study investigates the factors associated with the timing of corrective actions for retracted paper citations on English Wikipedia. We analyze various features, including a range of features related to the paper itself (e.g., its academic domain, open access status, and academic citation count). Additionally, we examine factors related to the retraction event (e.g. the reasons for retraction, time from publication to retraction, and editorial events before retraction). Finally, we analyze Wikipedia-specific factors such as RetractionBot flagging, page reference count, media count, categories count, and the level of editorial activity (e.g., the number of unique editors). These factors provide insight into the challenges of identifying and addressing problematic citations.
Our findings show that Wikipedia’s responsiveness to citations of retracted papers is inconsistent and often slow. While some citations are corrected quickly, many persist uncorrected for years. Among the 1,181 citation pairs analyzed, we found 71.6% represented problematic cases requiring repair: specifically, 51.5% were added before the retraction, and 20.1% were introduced afterwards without in-text warnings. The remaining 28.4% were classified as non-problematic because they explicitly noted the retraction status in-text at the time of citation. While many of these problematic citations are eventually corrected, our analysis reveals that they persist for a median of over 3.68 years (1,344 days). In the extended Cox model, the correction hazards are higher for citations that are flagged by RetractionBot and for citations that were framed as controversial at the time of addition. Higher correction hazards are also associated with open access papers and with greater Twitter/X attention prior to retraction, although the Twitter/X association is time-dependent and attenuates over time. Page-level structural integration is also associated with faster correction: pages with more categories show higher correction hazards. Conversely, a higher academic citation count is associated with a lower correction hazard, which may reflect the difficulty of challenging sources perceived as highly authoritative. We do not observe a stable average difference by retraction reason, though the association appears to vary over time. Together, these findings indicate that visibility cues and page organization are associated with the timing of citation-level repair on Wikipedia.
@@
Our paper contributes to CSCW research on collaborative knowledge infrastructures in four ways. First, it shifts the focus from Wikipedia’s well-documented responsiveness to high-salience events to its underexamined handling of low-visibility epistemic breakdowns, such as retracted scientific papers. Second, it provides a large-scale empirical analysis of how retracted citations persist—or are corrected—across academic domains, revealing how features such as social media signals, open access status, automated tools (e.g., RetractionBot), and page-level organization (e.g., number of categories) are associated with repair outcomes. Third, it highlights a critical division of labor between reference-level flagging and substantive in-text repair. While our results show that automated interventions (e.g., RetractionBot) significantly accelerate the correction process, they primarily serve as technical signals that catalyze human editorial work rather than replace it. This indicates a broader design gap between machine-generated alerts and editorial workflows: while machine-generated alerts are effective at surfacing issues, the contextual reader-facing correction remains a distinctly human task. Finally, it provides actionable opportunities for design, suggesting how Wikipedia’s tools, workflows, or interfaces might better support editors in identifying and repairing citation-level credibility failures.
@@ 2.1 Wikipedia as an Epistemic Infrastructure
Wikipedia has served as a prime example in CSCW and HCI research for studying large-scale, decentralized collaboration. As a social computing system, it depends on the interplay of human contributors, automated bots, and governance policies to produce and maintain public knowledge Forte and Lampe (2013); Halfaker and Riedl (2012); Geiger and Ribes (2010). This sociotechnical infrastructure supports a complex ecosystem of roles, routines, and technical tools that enable collaborative knowledge production Ackerman et al. (2013); Forte et al. (2012).
@@
Prior work has documented how Wikipedia adapts to rapidly changing external events. In crisis contexts such as natural disasters or breaking news, editors engage in high-tempo coordination, rapidly creating and revising articles to reflect emerging realities Keegan et al. (2013, 2012). Studies of social movement coverage, such as the documentation of Black Lives Matter, extend this perspective to the formation of collective memory in contentious and evolving domains Twyman et al. (2017). Keegan et al. Keegan and Brubaker (2015) show how editors update Wikipedia biographies after someone dies, working together to correct facts while also honoring the person’s memory. Recent research on the Russo-Ukrainian War extends this understanding to geopolitical conflicts, highlighting how editors mobilize to manage disruptive editing and maintain information integrity amidst information warfare Roberts and Xiong-Gum (2022); Kurek et al. (2025). Together, these studies underscore Wikipedia’s responsiveness, focusing on high-salience events that prompt rapid editorial attention.
@@
Yet recent work highlights Wikipedia’s vulnerabilities in less visible contexts. Kharazian et al. Kharazian et al. (2024) show how, in smaller language editions of Wikipedia, editorial control can become concentrated in the hands of a few users, allowing misinformation to persist. Similarly, Aragón et al. Aragón and Sáez-Trumper (2021) identify risks to knowledge integrity that can arise even without deliberate manipulation. Together, these studies show that Wikipedia’s openness—while foundational to its success—also means that no individual contributor is accountable for the accuracy of any specific claim. When responsibility is diffuse and editorial attention is uneven, epistemic breakdowns can persist without being noticed or repaired. Moreover, Wikipedia’s information credibility can face intentional challenges. Kumar et al. Kumar et al. (2016) studied the creation and detection of Wikipedia hoaxes, fabricated articles designed to deceive readers. Our study contributes to the literature by examining how Wikipedia handles a particular form of non-adversarial epistemic disruption: the citation of scientific articles that have been formally retracted. Failure to maintain the credibility of cited scientific sources leads to a non-negligible breakdown in information quality that is impactful regardless of intention.
@@
Prior work has provided important insights into Wikipedia’s collaborative maintenance processes at the article level. However, fewer studies have examined how the platform manages the credibility of cited knowledge when it becomes formally invalidated. Rather than tracing editorial workflows directly, we analyze the distribution of these citations across Wikipedia and identify factors associated with whether they are eventually corrected or persist unaddressed. This approach highlights how certain infrastructural and visibility features—such as automated flagging, attention signals, and page organization—are associated with whether citations are corrected or persist unaddressed.
@@ 2.2 Scientific Knowledge and Citations on Wikipedia
Scientific knowledge plays a central role in Wikipedia’s coverage of specialized topics, and recent scholarship has explored how scientific sources are cited, structured, and maintained. Prior work has mapped the landscape of scientific citations on Wikipedia, revealing both the breadth of disciplines represented and the dominant role of biomedical research and papers from high-impact journals. Singh, West, and Colavizza Singh et al. (2021) quantified this trend, finding that the life sciences and biomedicine dominate the cited literature, with journals such as Nature and Science among the most frequently referenced. Arroyo-Machado et al. Arroyo-Machado et al. (2020) similarly found that multidisciplinary and biomedical journals dominate Wikipedia’s scientific references. Nicholson et al. Nicholson et al. (2021) assessed citation quality, showing that Wikipedia tends to cite reputable journals, but still includes a notable fraction of sources with limited academic validation. Other studies have documented representational biases, such as the under-citation of women and non-Anglophone authors Zheng et al. (2023).
@@
Complementing this citation-level research, scholars have also examined how Wikipedia’s article quality ratings relate to citation behavior. Arroyo-Machado et al. Arroyo-Machado et al. (2022) found that articles rated as higher quality by the Wikipedia community tend to include more academic references, attract more edits, and receive higher page views—but that these associations vary depending on the topical domain. For example, the authors found that a high number of academic references strongly correlates with scholarly-oriented articles in fields like medicine and history. In contrast, high page views are more typical of articles about popular culture, and high talk page activity is associated with controversial subjects. More recently, Baigutanova et al. Baigutanova et al. (2023) conducted a large-scale longitudinal study of Wikipedia’s reference quality, finding that sourcing practices have generally improved over time, particularly with increasing reliance on peer-reviewed journals and scientific literature. This trend toward higher-quality sourcing builds on earlier observations about Wikipedia’s increasing preference for high-impact journals such as Nature and Science Nielsen (2007); Wedemeyer et al. (2008). Nevertheless, concerns persist about the reliability and maintenance of references, particularly in articles that receive limited editorial attention.
@@
While prior research has primarily focused on the inclusion and quality of citations, relatively little is known about how Wikipedia handles erroneous scientific citations, particularly those referencing retracted research. The study most closely aligned with ours is by Nicholson et al. Nicholson et al. (2021), who identified 118 retracted articles cited across 297 Wikipedia pages. In a manual sample of 50 citations, they found that about half failed to acknowledge the retraction, and even when retractions were mentioned, they were typically used illustratively rather than corrected or flagged clearly. Additionally, their study was based on a small sample. It did not examine what factors are associated with citation correction, thus creating an opportunity for a large-scale characterization of retraction persistence patterns across Wikipedia.
@@
We build on this line of work by identifying retracted papers cited on Wikipedia and analyzing how those citations are handled over time through article revisions. Our study offers a large-scale account of how retracted citations are noted, ignored, or removed. We also model the factors that are associated with these outcomes. In doing so, we extend prior research on citation practices by shifting attention from how scientific citations are sourced and evaluated to how they become invalid and are subsequently addressed. Our work also contributes to CSCW scholarship on collaborative maintenance and infrastructure resilience by examining how Wikipedia editors manage epistemic failures at the citation level.